how to calculate vector products like
(a*b)*(c*d) where * represent cross products

Grade:12

2 Answers

Himanshu
8 years ago
To calculate such product take a*b=m and then solve and later replace m by a*b 
A*b=m hence m*(c*d) solve by vector triple product
Ankit Jaiswal
8 years ago
 
To calculate such product take a*b=m and then solve and later replace m by a*b 
A*b=m hence m*(c*d) solve by vector triple product.